OpenStudy (anonymous):

Which is a unit vector? ⟨4953−−√,253−−√⟩ ⟨753−−√,453−−√⟩ ⟨753−−√,253−−√⟩ ⟨4953−−√,453−−√⟩

OpenStudy (anonymous):

\[<\frac{ 49 }{ \sqrt{53} },\frac{ \frac{ }{ 2 } }{ \sqrt{53} } \]

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

\[<\frac{ 49 }{ \sqrt{53} },\frac{ \frac{ }{ 2 } }{ \sqrt{53} } >\]

OpenStudy (anonymous):

I got the answer nvm

OpenStudy (freckles):

ok